OpenStudy (anonymous):

What is 5 pi over 3 radians in degrees?

OpenStudy (anonymous):

\[\frac{ degrees }{ 360 }=\frac{ 5π/3 }{ 2π }\] solve for degrees

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

the formula is to convert from degrees to radians and radians to degrees \[\frac{ degrees }{ 360 }=\frac{ radians }{ 2π }\] so plug where the radians is the 5π/3 and solve for degrees so you will get: \[degrees=\frac{ radians*360 }{ 2π }\] In your case : \[degrees=\frac{ 360\frac{ 5π }{ 3 } }{ 2π }=....\]

OpenStudy (anonymous):

300 degrees?

OpenStudy (anonymous):

yeah correct :)